The gas in a 250.0 mL piston experiences a change in pressure from 1.00 atm to 2.80 atm. What is the new volume (in mL) assuming the moles of gas and temperature are held constant?


1
Expert's answer
2021-02-10T01:46:18-0500

P1V1=P0V0

V1=P0V0/P1

V1=1.00*250.0/2.80=89.28(ml)